# PKS 2131--021 – Discovery of Strong Coherent Sinusoidal Variations from Radio to Optical Frequencies: Compelling Evidence for a Blazar Supermassive Black Hole Binary

### v1

@@ 6 The $\gamma$-ray Emission
Figure 10: Normalized OVRO and Fermi light curves. The red shaded areas indicate the extent of the increased $\gamma$-ray activity. The $\gamma$-ray data have been shifted in time by 638 days to align the peaks in the radio and $\gamma-$ray observations (see text).
@@
Although $\gamma$-rays show sharper flares compared to the broad radio peaks, it is clear that the enhanced $\gamma$-ray activity periods, when shifted, align very well with the radio flares (Fig. 10). Depending on the radio peak, we can achieve a good alignment between radio and $\gamma$-rays with time-lags ranging between 454 and 819 days. If we shift the $\gamma$-rays forward by 638 days to match the radio-optical phase lag we obtain the comparison shown in Fig. 10, which appears to us a reasonable match, given the uncertainties. This is consistent with the monotonic phase trend shown in Fig. 6(d), which includes the time-lag found between radio and the WISE observations of 470 days (see Section 4.2). The fact that the radio–$\gamma$-ray time-lags are consistent with what we find for the infrared and optical observations suggests inverse-Compton scattering for the $\gamma$-ray production mechanism (e.g., Liodakis et al., 2019).
7 Kinetic Orbital Model
@@
Figure 12: Schematic diagram (not to scale) showing the relationship between the typical locations of the emission at 345 GHz and 2.7 GHz on the KO model, for the case $\beta_{1}=0.020$, $\beta=0.995$, and $\gamma\sim 10$. The numbers in green script correspond to this case. This schematic illustrates the special case where the orbital plane and the jet axis are orthogonal.
@@
Suppose that we have a source at rest emitting isotropically and observed distantly with a flux density $S^{\prime}$. The observed flux density will be given by $S={\cal D}^{2-\alpha}S^{\prime}$ (Scheuer & Readhead, 1979), where $\alpha=d\ln S/d\ln\nu$ is the spectral index, and ${\cal D}$ is the Doppler factor.
@@
7.1 Consistency of the observed light curves with the Kinetic Orbital Model
In Appendix A, we show, based on MOJAVE VLBI measurements, that the Doppler factor of the core component in PKS 2131--021 is $D=14$, which, using the angle between the jet axis and the line of sight of $\theta=3.8^{\circ}$ (Homan et al., 2021), yields a Lorentz factor ${\rm\gamma}=10$ (see Appendix A). We can perform a simple consistency check with the KO model using the phase shifts we have determined. This also provides some useful constraints on the scales of the region producing the sinusoid in PKS 2131--021.
@@ 7.1 Consistency of the observed light curves with the Kinetic Orbital Model
$r^{\prime}>5\times 0.069\gamma^{2}{\rm pc}\sim 0.345\gamma^{2}\,{\rm pc}\,,$ (5)
so that, based on the value of $\gamma=10$,
$r^{\prime}\gtrsim 34.6{\rm pc,\,i.e.,\,\sim 113\,ly}.$ (6)
@@
$r^{\prime}_{\rm proj}=\theta r^{\prime}>0.345\gamma^{2}\theta\,{\rm pc}$ (7)
@@
$r^{\prime}_{\rm proj}\gtrsim 2.3{\rm pc,\,i.e.\,\sim 7.5\,ly},$ (8)
based on the values of $\theta=3.8^{\circ}$ derived by Homan et al. (2021) and $\gamma=10$.
@@
This fitting was done by eye, and it was done independently for the transitions at the end of the green sine wave and the beginning of the blue sine wave shown in Fig. 13. The mean levels without the sine waves, shown by the black horizontal lines differ by 2%, even though these fits were determined at transitions that are separated by 24 y. We also find that the mean level during the period when the sinusoidal variation was absent, marked by the black triangles, differs by only 7% from the mean of the residuals during the two periods when the sinusoidal variations were present. The good agreement between these mean levels lends credence to our hypothesis that there are underlying emission components that are not varying sinusoidally, shown by the black symbols in Fig. 13, and that the sinusoidal variations are seen as an addition to this underlying component.
@@
We have not attempted here to subtract the sinusoidal signal during the period from 2003 to 2008, when it appears to have been switching on, because the amplitude was clearly changing. Note that we the fractional variation in flux density of 0.72 is consistent with a value of $\beta_{1}=0.020$, and $\beta=0.995$, i.e. $\gamma=10$. For convenience we will adopt the value $\beta_{1}=0.020$.
@@
As an instructive exercise and guide, we consider the KO model of in the rest frame of the SMBHB, in the case where the orbital plane and the jet axis are orthogonal. In this case the locus of the jet is a helix on the surface of a cone, as shown in Fig. 12, and the opening angle of the cone is $2\beta_{1}/\beta$. From Equation 2 we see that with these parameters $\delta r^{\prime}$ between the regions emitting at 2.7 GHz and 345 GHz is 125.35 ly, and the distance along the jet axis between successive turns of the helix is 2.07 ly. Thus there are $\sim 60$ turns of the helix between these two surfaces. At first sight it may seem strange that the phase difference between these two emission regions separated by over 60 cycles is so small, but this is simply a result of the fact that the jet is advancing at nearly the speed of light ($0.995c$), so that for any distance light travels along the jet axis, the jet falls behind by only 1/200 of that distance, and thus, in $\sim$60 rotations the jet falls behind by only 0.3 of a cycle. The surface from which the emission emerges at a given frequency is at a fixed distance from the SMBH, but the jet is moving through this stationary surface at speed $v_{\textrm{jet}}$. This is indistinguishable from the situation considered by Rees (1967). Our conclusion for this section is that the data are consistent with the KO model.
@@
An important question is “At what level does one expect to see harmonics of the fundamental frequency?”. This is beyond the scope of the present paper, and will be discussed in a future paper, For the present, we refer the reader to Appendix B.
@@ 8 What Fraction of Blazars are SMBHB Candidates?
The recent discovery of a stochastic background of gravitational waves (GW) with periods of months to years (Agazie et al., 2023a; EPTA Collaboration et al., 2023) using millisecond pulsars as GW probes has promoted an intense search for the GW electromagnetic counterparts (e.g., Agazie et al., 2023b), which makes clear the importance of SMBHB searches such as the one we are engaged in with the OVRO 40 m Telescope. Millisecond pulsars (Readhead & Hewish, 1974; Backer et al., 1982) [^2] There were two crucial steps in the discovery of millisecond pulsars: (i) the discovery of interplanetary scintillation, at Galactic latitude $-0.3^{\circ}$, in 4C 21.53 by Readhead & Hewish (1974) (see Readhead (2024)), which first drew attention to the singular nature of this object; and (ii) the discovery of millisecond pulses from 4C 21.53W by Backer et al. (1982). enable searches for gravitational waves with periods of months to years using pulsar timing arrays such as the European pulsar timing array (EPTA, Antoniadis et al., 2023), the North American Nanohertz Observatory for Gravitational Waves (NANOGrav, Agazie et al., 2023a, c, b), and the Parkes pulsar timing array (PPTA, Zic et al., 2023). As it happens, PKS 2131--021 is almost optimally located for study with the EPTA, NANOGrav, and the PPTA. In Fig. 14, the black and white crosses mark PKS 2131--021 and the location of greatest NANOGrav sensitivity, respectively (Agazie et al., 2023b).
@@
Figure 14: Mollweide projection of the sky in right ascension and declination coordinates, showing the position of PKS 2131--021 (black cross) plotted on the NANOGrav strain upper limits (Agazie et al., 2023b). The gray curve shows the Galactic Plane. The white dot marks the location of the Galactic center. The white cross marks the position of the highest NANOGrav sensitivity. The red asterisks mark the positions of the millisecond pulsars used in the Agazie et al. (2023b) analysis, from which this figure is adapted. The black asterisk marks the position of 4C 21.53, the first millisecond pulsar.
@@
It is clearly important to determine the fraction of blazars that are SMBHB candidates. We are in the process of carrying out a detailed study applying the above GLS analysis to all of the AGN in the 40 m monitoring sample. Since this study is in progress we give here only a brief outline of our findings thus far. We have completed the local significance analysis on about half of the objects in the sample, and we find that about 15% of these pass the $3\sigma$ local significance threshold. We have yet to carry out the global significance analysis on these objects. In the case of PKS 2131--021 the ratio of the local $p$-value to the global $p$-value is $>$40. If we assume that the same applies to the other objects in the sample then we project that at most 20 of the objects in our sample will pass the $3\sigma$ global significance threshold. We have thus far found four AGN that pass the $3\sigma$ global significance threshold. Thus the projected fraction of AGN in our sample that are strong SMBHB candidates is at least 1 in 500, and could be 1 in 100.
@@
However, it should be borne in mind that the test we are applying is selecting SMBHB candidates in which the sinusoidal variation is both strong and persists over most of the 16 year period of observations, from 2008 to 2023, inclusive. We have seen in PKS 2131--021 that the sinusoidal variation was absent for 19 of the past 48 years. In addition the sinusoidal variations have to be strong enough, compared to the random variations in the source, to yield a $3\sigma$ global significance. It appears very likely that there will be AGN that are SMBHBs in which there are sinusodial signals that do not pass this threshold. For all of these reasons, in our view, the fraction of SMBHBs among jetted-AGN is likely to be $\sim 1$ in 100, or higher.
@@
It is interesting to note that JWST is finding a large number of dual AGN at high redshifts, some of which are separated by a few or several hundred parsecs (Perna et al., 2023; Marshall et al., 2023; Übler et al., 2023; Maiolino et al., 2023). The fraction of these dual AGN seems to be fairly high (20–30%), while the majority of AGN hosts have close companions or show post-merging signatures. While the fate of black holes in interacting galaxies at large (several-kiloprasec) separations is not yet clear (Di Matteo et al., 2023), black holes with separations of less than 1 kpc are expected to become black hole binaries on sinking times of a few to several hundred million years, especially given the high central stellar densities of these early galaxies, which increase the drag force on BHs. A fraction of such binary BHs with appropriate jet emission and orientation should be detectable as blazars, with periodic signals similar to those seen in PKS 2131--021 at $z=1.283$. It is difficult to compare the fraction of blazars that are SMBHB candidates and the fraction of candidate BH mergers from recent JWST observations with theoretical models and simulations of BH merging, due to poor statistics and large uncertainties in both observations and simulations. However, these are all areas of active and rapid development, and a quantitative assessment might be possible in the near future.
@@ 9 Conclusion and Future Work
    The periodicity at radio wavelengths has remained persistent and stable across several epochs. The newly discovered periodicity in optical and possibly in $\gamma$-rays again shows the robustness of this discovery. We thus predict that this periodicity will persist with additional radio through $\gamma$-ray monitoring of this source.

### v2

@@ 8 The $\gamma$-ray Emission
Figure 10: Normalized OVRO and Fermi light curves. The red shaded areas indicate the extent of the increased $\gamma$-ray activity. The $\gamma$-ray data have been shifted in time by 612 days to align the peaks in the radio and $\gamma$-ray observations (see text).
Although $\gamma$-rays show sharper flares compared to the broad radio peaks, it is clear that the enhanced $\gamma$-ray activity periods, when shifted, align very well with the radio flares (Fig. 10). Depending on the radio peak, we can achieve a good alignment between radio and $\gamma$-rays with time-lags ranging between 454 and 819 days. If we shift the $\gamma$-rays forward by 612 days to match the radio-optical phase lag we obtain the comparison shown in Fig. 10, which appears to us a reasonable match, given the uncertainties. This is consistent with the monotonic phase trend shown in Fig. 3(d), which includes the time-lag found between radio and the WISE observations of 470 days (see Section 6.1). The fact that the radio–$\gamma$-ray time-lags are consistent with what we find for the infrared and optical observations suggests inverse-Compton scattering for the $\gamma$-ray production mechanism (e.g., Liodakis et al., 2019).
9 Kinematic Orbital Model
@@
Figure 12: Schematic diagram (not to scale) showing the relationship between the typical locations of the emission at 345 GHz and 2.7 GHz on the KO model, for the case $\beta_{1}=0.020$, $\beta=0.995$, and $\gamma\sim 10$. The numbers in green script correspond to this case. On this model we assume that the dominant emission at each frequency arises from the region where the jet becomes optically thick at that frequency. This schematic illustrates the special case where the orbital plane and the jet axis are orthogonal, ($j=\pi/2$).
@@
Suppose we have a source emitting isotropically in its rest frame with a flux density $S^{\prime}$. The observed flux density will be given by $S={\cal D}^{2-\alpha}S^{\prime}$ (Scheuer & Readhead, 1979), where $\alpha=d\ln S/d\ln\nu$ is the spectral index, and ${\cal D}$ is the Doppler factor.
@@
9.1 Consistency of the observed light curves with the Kinematic Orbital Model
In Appendix C, we show, based on MOJAVE VLBI measurements, that the Doppler factor of the core component in PKS 2131--021 is $\cal{D}=14$, which, using the angle between the jet axis and the line of sight of $\theta=3.8^{\circ}$ (Homan et al., 2021), yields a Lorentz factor ${\rm\gamma}=10$ (see Appendix C). We can perform a simple consistency check with the KO model using the phase shifts we have determined. This also provides some useful constraints on the scales of the region producing the sinusoid in PKS 2131--021.
@@ 9.1 Consistency of the observed light curves with the Kinematic Orbital Model
$r^{\prime}>5\times 0.064\gamma^{2}{\rm pc}\sim 0.320\gamma^{2}\,{\rm pc}\,,$ (6)
so that, based on the value of $\gamma=10.0125\;(\beta=0.995)$,
$r^{\prime}\gtrsim 32.1\,{\rm pc,\,i.e.,\,\sim 105\,ly}.$ (7)
@@
$r^{\prime}_{\rm proj}=\theta r^{\prime}>0.320\gamma^{2}\theta\,{\rm pc}$ (8)
@@
[…]
based on the value of $\theta=3.8^{\circ}$ derived by Homan et al. (2021).
@@
This fitting was done by eye, and it was done independently for the transitions at the end of the green sine wave and the beginning of the blue sine wave shown in Fig. 13. The mean levels without the sine waves, shown by the black horizontal lines differ by 2%, even though these fits were determined at transitions that are separated by 24 yr. We also find that the mean level during the period when the sinusoidal variation was absent, marked by the black triangles, differs by only 7% from the mean of the residuals during the two periods when the sinusoidal variations were present. The good agreement between these mean levels lends credence to our hypothesis that there are underlying emission components that are not varying sinusoidally, represented by the black symbols in Fig. 13, and that the sinusoidal variations are seen as an addition to this underlying component.
@@
We have not attempted here to subtract the sinusoidal signal during the period from 2003 to 2008, when it appears to have been switching on, because the amplitude was clearly changing. Note that we the fractional variation in flux density of 0.72 is consistent with a value of $\beta_{1}=0.020$, and $\beta=0.995$, i.e. $\gamma\sim 10$. For convenience we will adopt the value $\beta_{1}=0.020$.
@@
As an instructive exercise and guide, we consider the KO model in the rest frame of the SMBHB, in the case where the orbital plane and the jet axis are orthogonal. In this case the locus of the jet is a helix on the surface of a cone, as shown in Fig. 12, and the opening angle of the cone is $2\beta_{1}/\beta$. From Equation 3 we see that with these parameters $\delta r^{\prime}$ between the regions emitting at 2.7 GHz and 345 GHz is 116.32 ly, and the distance along the jet axis between successive turns of the helix is 2.07 ly. Thus there are $\sim 60$ turns of the helix between these two surfaces. At first sight it may seem strange that the phase difference between these two emission regions separated by over 60 cycles is so small, but this is simply a result of the fact that the jet is advancing at nearly the speed of light ($0.995c$), so that for any distance light travels along the jet axis, the jet falls behind by only 1/200 of that distance, and thus, in $\sim$60 rotations the jet falls behind by only 0.3 of a cycle. The surface from which the emission emerges at a given frequency is at a fixed distance from the SMBH, but the jet is moving through this stationary surface at speed $v_{\textrm{jet}}$. This is indistinguishable from the situation considered by Rees (1967). Our conclusion for this section is that the data are consistent with the KO model. It should be clear from Fig. 12 that the phase coherence we observe in PKS 2131--021 across a wide range of observing frequencies is due to the fact that $\theta\ll 1$. Were this not the case, the light travel time arguments from the emission regions to the observer illustrated in Fig. 12 would not apply and coherence between different frequency bands would be lost.
@@
An important question is “At what level does one expect to see harmonics of the fundamental frequency?”. This is beyond the scope of the present paper, and will be discussed in a future paper, For the present, we refer the reader to Appendix D.
@@ 10 What Fraction of Blazars are SMBHB Candidates?
The recent discovery of a stochastic background of gravitational waves (GW) with periods of months to years (Agazie et al., 2023a; EPTA Collaboration et al., 2023) using millisecond pulsars as GW probes has promoted an intense search for the GW electromagnetic counterparts (e.g., Agazie et al., 2023b), which makes clear the importance of SMBHB searches such as the one we are engaged in with the OVRO 40 m Telescope. Millisecond pulsars (Readhead & Hewish, 1974; Backer et al., 1982) [^2] There were two crucial steps in the discovery of millisecond pulsars: (i) the discovery of interplanetary scintillation, at Galactic latitude $-0.3^{\circ}$, in 4C 21.53 by Readhead & Hewish (1974) (see Readhead (2024)), which first drew attention to the singular nature of this object; and (ii) the discovery of millisecond pulses from 4C 21.53W by Backer et al. (1982). enable searches for gravitational waves with periods of months to years using pulsar timing arrays such as the European pulsar timing array (EPTA, Antoniadis et al., 2023), the North American Nanohertz Observatory for Gravitational Waves (NANOGrav, Agazie et al., 2023a, c, b), the Parkes pulsar timing array (PPTA, Zic et al., 2023), and the MeerKAT Pulsar Timing Array (MPTA, Miles et al., 2025). As it happens, PKS 2131--021 is almost optimally located for study with the EPTA, NANOGrav, and the PPTA. In Fig. 14, the black and white crosses mark PKS 2131--021 and the location of greatest NANOGrav sensitivity, respectively (Agazie et al., 2023b).
@@
Figure 14: Mollweide projection of the sky in right ascension and declination coordinates, showing the position of PKS 2131--021 (black cross) plotted on the NANOGrav strain upper limits (Agazie et al., 2023b). The gray curve shows the Galactic Plane. The white dot marks the location of the Galactic center. The white cross marks the position of the highest NANOGrav sensitivity. The red asterisks mark the positions of the millisecond pulsars used in the Agazie et al. (2023b) analysis, from which this figure is adapted. The black asterisk marks the position of 4C 21.53, the first millisecond pulsar (see text).
It is clearly important to determine the fraction of blazars that are strong SMBHB candidates. Sandrinelli et al. (2017, 2018) estimated that $\sim 10\%$ of $\gamma$-ray bright blazars are QPO candidates, but they later determined that all of their candidates were likely produced by the red noise tail (Covino et al., 2019). Ackermann et al. (2015) reported QPOs in PG 1553+113 based on the Fermi-LAT $\gamma$-ray light curve over a period of 6.9 years. This periodicity has recently been confirmed using 15 years of Fermi-LAT data (Abdollahi et al., 2024), see also Aniello et al. (2024). In a multi-wavelength study MAGIC Collaboration et al. (2024) used the above GLS analysis (with 20,000 simulations) to show that the QPOs in PG 1553+113 have a global $p$-value $=1\times 10^{-3}\;(3.2\sigma)$ in the Fermi light curve, and $p$-value $=3.4\times 10^{-2}$ in the OVRO 15 GHz light curve. Thus PG 1553+113 passes the $3\sigma$ global $p$-value test at $\gamma$-ray energies and is a strong SMBHB candidate, according to the criterion we adopted in Paper 1.
de la Parra et al. (2024) have similarly shown that PKS J0805--0111 is a strong SMBHB candidate, based on the OVRO 15 GHz light curve. They also found hints of the same periodicity in the Fermi-LAT light curve.
Thus, in the complete statistical sample of 1158 blazars being monitored at OVRO (Richards et al., 2011), three strong SMBHB candidates have been identified thus far: one of them (PG 1553+113) exhibits clear periodic oscillations at $\gamma$-ray energies and hints of periodic oscillations at radio energies; while the other two (PKS J0805--0111 and PKS 2131--021) exhibit clear periodic oscillations at radio energies and hints of periodic oscillations at $\gamma$-ray energies. It is clear, therefore, that the fraction of blazars that are strong SMBHB candidates, with periods in the range of a few months to $\sim$ five years, is $\gtrsim 1$ in 400.
@@
We are in the process of carrying out a detailed study applying the above GLS analysis to all of the $\sim 1800$ AGN in the 40 m monitoring sample. Since this study is in progress we give here only a brief outline of our findings thus far. We have completed the local significance analysis on about half of the objects in the sample, and we find that about 15% of these pass the $3\sigma$ local significance threshold. We have yet to carry out the global significance analysis on these objects. In the case of PKS 2131--021 the ratio of the local $p$-value to the global $p$-value is $>$40. If we assume that the same applies to the other objects in the sample then we project that at most 20 of the objects in our sample will pass the $3\sigma$ global significance threshold. We have thus far found four AGN that pass the $3\sigma$ global significance threshold. Thus the projected fraction of AGN in our sample that are strong SMBHB candidates, based on the radio data alone, is at least 1 in 500.
@@
Thus far we have picked out only the strongest SMBHB candidates - i.e. those that show very clear periodicity at $\gamma$-ray or radio frequencies. Clearly, there must be other SMBHB candidates in which the periodicities do not stand out so clearly against the background of strong variability we see in all blazars. It should be borne in mind that the test we are applying to the radio light curves is selecting SMBHB candidates in which the sinusoidal variation is both strong and persists over most of the 16 year period of observations, from 2008 to 2023, inclusive. We have seen in PKS 2131--021 that the sinusoidal variation was absent for 19 of the past 48 years. In addition the sinusoidal variations have to be strong enough, compared to the random variations in the source, to yield a $3\sigma$ global significance. It appears very likely that there will be AGN that are SMBHBs in which there are sinusodial signals that do not pass this threshold. For all of these reasons, in our view, the fraction of SMBHBs among jetted-AGN is likely to be $\sim 1$ in 100, or higher.
@@
It is interesting to note that JWST is finding a large number of dual AGN at high redshifts, some of which are separated by a few or several hundred parsecs (Perna et al., 2023; Marshall et al., 2023; Übler et al., 2023; Maiolino et al., 2023). The fraction of these dual AGN seems to be fairly high (20–30%), while the majority of AGN hosts have close companions or show post-merging signatures. While the fate of black holes in interacting galaxies at large (several-kiloparsec) separations is not yet clear (Di Matteo et al., 2023), black holes with separations of less than 1 kpc are expected to become black hole binaries on sinking times of a few to several hundred million years, especially given the high central stellar densities of these early galaxies, which increase the drag force on BHs. A fraction of such binary BHs with appropriate jet emission and orientation should be detectable as blazars, with periodic signals similar to those seen in PKS 2131--021 at $z=1.283$. It is difficult to compare the fraction of blazars that are SMBHB candidates and the fraction of candidate BH mergers from recent JWST observations with theoretical models and simulations of BH merging, due to poor statistics and large uncertainties in both observations and simulations. However, these are all areas of active and rapid development, and a quantitative assessment might be possible in the near future.
@@ 11 Conclusion and Future Work
    The periodicity at radio wavelengths has remained persistent and stable across several epochs. The newly discovered periodicity in optical and possibly in $\gamma$-rays again shows the robustness of this discovery. We thus predict that this periodicity will continue to be observed in this object over a wide range of frequencies.
@@
An important next step in exploring sources like PKS 2131--021 is to seek spectroscopic manifestation of the periodicity. This is an approach which has been highly productive in many Galactic binary sources, most notably SS 433 (see, e.g., Schmidtobreick & Blundell, 2006) which exhibits “stationary” and “moving” spectral lines plus strong evidence for a third body, all of which exhibit many periodic behaviors reflecting an underlying, dynamical clock.
There are many ways in which variable emission lines might arise in sources such as PKS 2131--021. These include launching a high speed wind from an orbiting accretion disk, either radiatively or magnetically; forming lines in an expanding and cooling high-speed jet which might orbit and/or precess and or be perturbed by the companion black hole; and having the proposed black hole binary being accompanied by stars or even a third black hole. Indeed, it would be rather surprising if the various dynamical explanations, that we have entertained to account for the periodicity in the flux, produced no associated spectroscopic variation.
[…]
If the Kinetic Orbital (KO) model is correct, then the optical sinusoidal curves will give us the expected phase of the GW waves from the SMBHB candidates. At this point only the expected amplitude of the GW will be unknown. Clearly, were a method to be developed for measuring the speeds of the SMBHBs, then a fully-coherent search for GW from the combination of ${\sim}100$ SMBHB candidates could be mounted in which the period, the phase, and the amplitude of the expected GW signal from each candidate could be added coherently. In such circumstances, the chances of being able to confirm, the SMBHB origin of the stochastic background of GW would be greatly increased.